OpenStudy (secret-ninja):

Solve for x: -3|x + 7| = -12 x = 5 over 3, x = −19 over 3 x = −3, x = −11 x = −3, x = 11 No solution

OpenStudy (haseeb96):

just repeat that my previous method again and get the answer

OpenStudy (texaschic101):

-3 | x + 7 | = -12 --- divide both sides by -3 | x + 7 | = 4 (x + 7) = 4 --- now solve for x -(x + 7) = 4 -- now solve for x the answers you get should match an answer choice

OpenStudy (secret-ninja):

So B? lol

OpenStudy (haseeb96):

yes